# TableRAG: A Retrieval Augmented Generation Framework for Heterogeneous Document Reasoning

### v1

@@ Abstract
Retrieval-Augmented Generation (RAG) has demonstrated considerable effectiveness in open-domain question answering. However, when applied to heterogeneous documents, comprising both textual and tabular components, existing RAG approaches exhibit critical limitations. The prevailing practice of flattening tables and chunking strategies disrupts the intrinsic tabular structure, leads to information loss, and undermines the reasoning capabilities of LLMs in multi-hop, global queries. To address these challenges, we propose TableRAG, an hybrid framework that unifies textual understanding and complex manipulations over tabular data. TableRAG iteratively operates in four steps: context-sensitive query decomposition, text retrieval, SQL programming and execution, and compositional intermediate answer generation. We also develop HeteQA, a novel benchmark designed to evaluate the multi-hop heterogeneous reasoning capabilities. Experimental results demonstrate that TableRAG consistently outperforms existing baselines on both public datasets and our HeteQA, establishing a new state-of-the-art for heterogeneous document question answering. We release TableRAG at https://github.com/yxh-y/TableRAG/tree/main.
@@
Xiaohan Yu∗ Huawei Cloud BU, Beijing yuxiaohan5@huawei.com Pu Jian∗ Huawei Cloud BU, Beijing jianpu2@huawei.com Chong Chen† Huawei Cloud BU, Beijing chenchong55@huawei.com
@@ 1 Introduction
To address these limitations of existing RAG systems, we propose TableRAG, an hybrid (SQL execution and text retrieval) framework that dynamic transitions between textual understanding and complex manipulations over tabular data. TableRAG interacts with tables by leveraging SQL as an interface. Concretely, the framework operates via a two-stage process: an offline database construction phase and an online inference phase of iterative reasoning. The iterative reasoning procedure comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) SQL programming and execution, and (iv) intermediate answer generation. The utilization of SQL enables precise symbolic execution by treating table-related queries as indivisible reasoning units, thereby enhancing both computational efficiency and reasoning fidelity. To facilitate rigorous evaluation of multi-hop reasoning over heterogeneous documents, we introduce HeteQA, a novel benchmark consisting of 304 examples across nine diverse domains. Each example contains a composition across five distinct tabular operations. We evaluate TableRAG on both established public benchmarks and our HeteQA dataset against strong baselines, including generic RAG and program-aided approaches. Experimental results demonstrate that TableRAG consistently achieves state-of-the-art performance. Overall, our contributions are summarized as follows:
@@
We propose TableRAG, an hybrid framework that unifies textual understanding and complex manipulations over tabular data. TableRAG comprises an offline database construction phase and a four-step online iterative reasoning process.
@@ 3.1 Overview and Design Principles
We propose TableRAG, an hybrid (SQL execution and text retrieval) framework designed to preserve table structural integrity and facilitate heterogeneous reasoning. As depicted in Figure 2, TableRAG consists of offline and online workflows. The offline phase is tasked with database construction, while the online phase facilitates iterative reasoning. The reasoning procedure unfolds in a four-stage process: (i) Context-sensitive query decomposition, which identifies the respective roles of textual and tabular modalities within the query. (ii) Text retrieval. (iii) SQL programming and execution, which is selectively invoked for subqueries requiring tabular data reasoning. (iv) Compositional intermediate answer generation. The preferential use of SQL is motivated by its capacity to leverage the expressive strength of symbolic execution over structured data, thereby enabling tabular components within user queries to be treated as monolithic reasoning units. In contrast, other languages like Python incur substantial computational overhead when dealing with large-scale data or complex workloads (Shahrokhi et al., 2024).
@@ 3.3 Iterative Reasoning
To address multi-hop, global queries that require compositional reasoning over texts and tables, we introduce an iterative inference process aligned with in Equation 1. This process comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) program and execute SQL, and (iv) compositional intermediate answer generation. Through repeated cycles of decomposition and resolution, a solution to the query is progressively constructed. Detailed prompt templates are provided in Appendix F.
@@ 4.1 Data Collection
To support queries that integrate both tabular and textual information, we augment the instance by leveraging the associated Wikipedia document. Specifically, certain entities within the query are replaced with reference-based formulations by the human annotators. For example, the query "Which driver " can be rephrased as "What is the nationality of the driver ". This entity substitution can either modify the subject of the question and its corresponding answer or alter the query phrasing while preserving the original answer. The annotation guidelines and annotator profiles are detailed in Appendix B.
@@ 5.1.2 Implementation Details
In the text retrieval process, we employ the BGE-M3 series models (Chen et al., 2024a, b). During recall, we retain the top 30 candidates, from which the top 3 are subsequently selected via reranking. To manage large inputs, the text is chunked into segments of 1000 tokens, with a 200-token overlap between consecutive chunks. The iterative loop is bounded by a maximum of 5 iterations. For backbone LLMs, we utilize Claude-3.5-Sonnet as a representative closed-source LLM, while Deepseek-V3, Deepseek-R1 (Guo et al., 2025), and Qwen-2.5-72B (Yang et al., 2024) serve as the open-source counterparts. A consistent backend is maintained for all modules in the online iterative reasoning process. We use accuracy as the evaluation metric, assessed by Qwen-2.5-72B, which yields a binary score of 0 or 1. The prompt is shown in Appendix F.
@@ 5.1.3 Baselines
We evaluate the performance of TableRAG by benchmarking it against three distinct baseline methodologies: (1) Direct answer generation with LLMs. (2) NaiveRAG, which processes tabular data as linearized Markdown formatted texts and subsequently applies a standard RAG pipeline. (3) React (Yao et al., 2023), a prompt based paradigm to synergize reasoning and acting in LLMs with external knowledge sources. (4) TableGPT2 (Su et al., 2024) employs a Python-based execution module to generate code (e.g., Pandas) for answer derivation within a simulated environment. The detailed implementation of these baseline methods is provided in Appendix C.
@@ 5.2 Main Result
The main results across different LLMs backbones are presented in Table 1. Several key observations emerge: (1) The ReAct framework demonstrates advantages over naive RAG on multi-source data, but exhibits degraded performance on single-source data that requires tabular reasoning. This can be attributed to context insensitivity during multi-turn reasoning. Queries decomposed into multiple sub-tasks, such as filtering or aggregation, suffering from information incompleteness or error propagation. (2) TableGPT2 yields acceptable results solely on single-source queries, such as WikiTQ, underscoring its limited capacity for handling multi-source queries. This reflects a lack of generalizability in heterogeneous information environments. (3) TableRAG surpasses all baselines, achieving at least a 10% improvement over the strongest alternative. Notably, it performs robustly across both single-source and multi-source data. This performance gain is attributed to the incorporation of symbolic reasoning, which enables effective adaptation to heterogeneous documents. Moreover, the consistency in performance across different LLM backbones underscores TableRAG’s architectural generality and compatibility with a broad range of backbones.
@@ 7 Analysis
We provide a comprehensive analysis of TableRAG in this section, with additional results presented in Appendix D.
@@ Appendix A HeteQA
A.1 Prompt Design
The templates employed for HeteQA benchmark construction are presented in Table LABEL:SQL.
A.2 Table Collection
To enable complex reasoning over tabular structures, we curate a subset of extensive tables from the Wikipedia-based corpus. Specifically, we retain only tables containing more than 20 rows and at least 7 columns. This filtering process reduced the initial collection from 15,314 tables to 1,345 candidates that meet the criteria for structural richness. To further enhance query diversity and reduce redundancy in the dataset, we apply a deduplication step based on schema similarity. In cases where multiple tables share identical column structures — for example, the entries for 1947 BAA draft and 1949 BAA draft — only a single representative instance is preserved. This procedure yielded a final dataset comprising 155 unique tables.
A.3 Data Collection
As illustrated in Figure 8, each data instance is constructed through a three-stage pipeline, facilitated via the Claude 3.7 Sonnet API333https://www.anthropic.com/claude/sonnet.
To construct the dataset, we prompt the LLM to generate SQL or pandas code associated with queries conditioned on a provided table schema. The generated code is then executed against the table to obtain the corresponding answer. Instances that fail to produce an executable result are discarded. To promote diversity within the dataset and avoid redundancy, we eliminate any instances exhibiting duplicate queries or identical answers. This filtering is performed using regular expressions to detect lexical or semantic repetition. Additionally, we discard cases where the execution result is ambiguous, such as queries seeking the top-ranked entity when multiple candidates are tied. Following automated filtering, all remaining instances undergo a manual verification process. Two human annotators independently assess each query, the corresponding code, and the resulting answer. In cases of inconsistency or error, annotators revise the SQL or Python code and correct the associated answer to ensure accuracy and coherence.
Appendix B Guidelines for HeteQA Annotations
B.1 Annotator Profiles
The two annotators are volunteers whose native language is Chinese and who possess fluent English proficiency. Both annotators are professional software engineers with substantial experience in SQL and Python programming.
B.2 Guidelines for Answer Verification
Objective
Your task is to verify whether the provided answer correctly and completely addresses the natural language query, based on the associated code and table. You will also identify and handle ambiguous or non-unique cases.
Annotation Steps
1. Verify the Answer
Read the natural language query and examine the accompanying Excel table, SQL statements, and Python (pandas) code. Verify whether the output aligns with the semantics of the original query and the underlying data. To this end, you are permitted to execute multiple operations on the provided Excel files.
2. Correction Tasks (if necessary)
If the answer is incorrect, modify both the code and the answer so that they correctly fulfill the query. Ensure the modified code is minimal, clean, and logically sound.
3. Ambiguity and Tie Cases
In instances where the query yields no unique resolution - whether due to tied outcomes, under specified conditions, or semantic ambiguity in the formulation - you can employ principled strategies to ensure meaningful processing:
- Option A: Modify the query to resolve the ambiguity (e.g., make clarification).
- Option B: If the ambiguity is irreparable or the answer depends on arbitrary choices, discard the case.
Additional Guidelines
Maintain consistency between the query, code, and answer. Ensure your corrections do not introduce new ambiguities or assumptions not grounded in the table or query.
B.3 Guidelines for Document Reference
Objective
Your task is to add an additional reasoning hop to the original query using the provided Wikipedia entities and content. This will help transform the original query into a more complex, multi-hop question that requires deeper reasoning.
Input Data
Each data case consists of:
Original Query
Answer
A set of Wikipedia entities relevant to the query or the answer
Corresponding Wikipedia content for each entity
Task Overview
For each data case, check if the original answer is mentioned in the provided Wiki entities:
- If not mentioned or ambiguous (e.g., "Jordan" for both basketball star and sports brand), do not change the query and answer. Just mark the case as “No modification needed”.
- If the answer entity exists in the Wikipedia content, perform the following steps:
1. Identify key factual descriptions about the answer entity from its Wiki content.
2. Add a reasoning hop to the original query that leads to the answer via this key fact.
3. Generate two candidate (Query, Answer) pairs by rewriting the query to incorporate this extra reasoning step and updating the answer accordingly.
Example
Original Query: Which album takes first place on the Billboard leaderboard in 2013?
Original Answer: ArtPop
Wiki Entity: ArtPop (album)
A key description: It was released on November 6, 2013, by Streamline and Interscope Records.
Modified Query Candidates:
Who released the album that takes the first place on the Billboard leaderboard in 2013?
→ Answer: Streamline and Interscope Records
Appendix C Implementation Details of Baselines
This section provides a detailed account of the implementation procedures for the baseline methodologies to ensure a fair and reproducible evaluation.
C.1 ReAct
For the ReAct framework, we preprocess tabular data by converting it into markdown-formatted plain text. To ensure consistency in experimental conditions, we adopt the same chunking and retrieval configurations as those employed in our TableRAG model. We build upon the publicly available ReAct implementation444https://github.com/ysymyth/ReAct. The framework addresses user queries through an iterative reasoning loop consisting of Thought, Action, and Observation steps, culminating in a final answer via the Finish operation. The max iteration is set to 5, the same as TableRAG.
C.2 TableGPT2
We evaluate TableGPT2 using its officially released TableGPT Agent implementation 555https://github.com/tablegpt/tablegpt-agent. As specified in its API documentation, we provide both the document content and the tabular data as inputs to the HumanMessage class, ensuring adherence to the intended usage of the model:
@@ C.2 TableGPT2
[…]
| sql_ans | Kath & Kimderella |
| answer | Riley, Turner, and Magda Szubanski |
Appendix D Analysis
We present a series of auxiliary analyses conducted on both our proposed benchmark, HeteQA, and the publicly available HybridQA dataset. These analyses offer further insight into the behavior and limitations of TableRAG beyond the primary evaluation metrics.
D.1 Analysis on HybridQA
We extend our investigation to the HybridQA dataset by examining the performance distribution on data domains and conducting a detailed error analysis, using DeepSeek-v3 as the backbone model. The results, summarized in Figures 10 and 9, reveal patterns consistent with those observed on HeteQA. This parallel further validates the generality of our observations across HeteQA settings.
D.2 HyperParameter Analysis
To elucidate the influence of the top- retrieval parameter on the performance of TableRAG, we undertook a systematic sensitivity analysis. While our main experimental setup fixed , we expanded our investigation on the HeteQA dataset, utilizing DeepSeek-V3 as the retrieval backbone, and varied across the set . The resulting performance metrics are illustrated in Figure 11.
Our observations reveal distinct behaviors across benchmarks with respect to the choice of . Notably, HybridQA exhibits superior performance at higher values. This effect is plausibly attributable to the lack of deduplication within its tables, whereby overlapping or similar information from multiple tables or Wikipedia documents contributes constructively during retrieval. In aggregate, setting strikes an effective balance, yielding robust performance across both benchmarks while maintaining computational efficiency. This efficiency gain arises from the direct relationship between the top- retrieval size and the subsequent LLM context length, underscoring the practical importance of this hyperparameter in optimizing the trade-off between accuracy and resource consumption.
Appendix E Check List
Harmful information And Privacy
We propose a new RAG solution to address multi-hop problems related to heterogeneous data, without involving any harmful information or privacy. In addition, we provide a heterogeneous benchmark containing tables and texts sourced from Wikipedia. All of the data is publicly available, contains no personal information, and involves no harmful content.
License and Intend
We provide the license we used here:
Claude 3.5 Sonnet (https://www.anthropic.com/legal/aup)
Qwen2.5-72B-Instruct (https://huggingface.co/Qwen/Qwen2.5-72B-Instruct/blob/main/LICENSE)
TableGPT Agent (Apache License 2.0) and TableGPT2-7B (https://huggingface.co/tablegpt/TableGPT2-7B/blob/main/LICENSE)
Our use of these existing artifacts was consistent with their intended use.
Documentation of the artifacts
We propose a novel Retrieval-Augmented Generation (RAG) framework that integrates traditional document retrieval with structured data querying via SQL, aiming to enhance performance in table-related question answering tasks. The framework comprises two fundamental stages: offline database construction and online interactive reasoning. Compared to conventional document-centric RAG approaches, our architecture offers additional reliable SQL execution results—when table fragments are involved in the retrieval process. This supplementary source mitigates the limitations of generative models in handling structured tabular data.
To facilitate a more comprehensive evaluation of our framework, we further construct a heterogeneous benchmark named HeteQA. This benchmark aims to evaluate the capability to handle multi-hop reasoning tasks across heterogeneous documents. The benchmark instances are initially generated by LLMs, and then rigorously validated by experienced programmers and database administrators to ensure correctness and realism.
Appendix F Prompts
The prompts for TableRAG are presented in this section.

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) has demonstrated considerable effectiveness in open-domain question answering. However, when applied to heterogeneous documents, comprising both textual and tabular components, existing RAG approaches exhibit critical limitations. The prevailing practice of flattening tables and chunking strategies disrupts the intrinsic tabular structure, leads to information loss, and undermines the reasoning capabilities of LLMs in multi-hop, global queries. To address these challenges, we propose TableRAG, an SQL-based framework that unifies textual understanding and complex manipulations over tabular data. TableRAG iteratively operates in four steps: context-sensitive query decomposition, text retrieval, SQL programming and execution, and compositional intermediate answer generation. We also develop HeteQA, a novel benchmark designed to evaluate the multi-hop heterogeneous reasoning capabilities. Experimental results demonstrate that TableRAG consistently outperforms existing baselines on both public datasets and our HeteQA, establishing a new state-of-the-art for heterogeneous document question answering.
@@
Xiaohan Yu∗, Pu Jian∗, Chong Chen#
Huawei Cloud BU, Beijing
{yuxiaohan5, jinapu2, chenchong55}@huawei.com
https://github.com/yxh-y/TableRAG
@@ 1 Introduction
To address these limitations of existing RAG systems, we propose TableRAG, an SQL-based framework that dynamically transitions between textual understanding and complex manipulations over tabular data. TableRAG interacts with tables by leveraging SQL as an interface. Concretely, the framework operates via a two-stage process: an offline database construction phase and an online inference phase of iterative reasoning. The iterative reasoning procedure comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) SQL programming and execution, and (iv) intermediate answer generation. The utilization of SQL enables precise symbolic execution by treating table-related queries as indivisible reasoning units, thereby enhancing both computational efficiency and reasoning fidelity. To facilitate rigorous evaluation of multi-hop reasoning over heterogeneous documents, we introduce HeteQA, a novel benchmark consisting of 304 examples across nine diverse domains. Each example contains a composition across five distinct tabular operations. We evaluate TableRAG on both established public benchmarks and our HeteQA dataset against strong baselines, including generic RAG and program-aided approaches. Experimental results demonstrate that TableRAG consistently achieves state-of-the-art performance. Overall, our contributions are summarized as follows:
@@
We propose TableRAG, an SQL-based framework that unifies textual understanding and complex manipulations over tabular data. TableRAG comprises an offline database construction phase and a four-step online iterative reasoning process.
@@ 3.1 Overview and Design Principles
We propose TableRAG, an SQL-based framework designed to preserve table structural integrity and facilitate heterogeneous reasoning. As depicted in Figure 2, TableRAG consists of offline and online workflows. The offline phase is tasked with database construction, while the online phase facilitates iterative reasoning. The reasoning procedure unfolds in a four-stage process: (i) Context-sensitive query decomposition, which identifies the respective roles of textual and tabular modalities within the query. (ii) Text retrieval. (iii) SQL programming and execution, which is selectively invoked for subqueries requiring tabular data reasoning. (iv) Compositional intermediate answer generation. The preferential use of SQL is motivated by its capacity to leverage the expressive strength of symbolic execution over structured data, thereby enabling tabular components within user queries to be treated as monolithic reasoning units. In contrast, other languages like Python incur substantial computational overhead when dealing with large-scale data or complex workloads (Shahrokhi et al., 2024).
@@ 3.3 Iterative Reasoning
To address multi-hop, global queries that require compositional reasoning over texts and tables, we introduce an iterative inference process aligned with in Equation 1. This process comprises four core operations: (i) context-sensitive query decomposition, (ii) text retrieval, (iii) program and execute SQL, and (iv) compositional intermediate answer generation. Through repeated cycles of decomposition and resolution, a solution to the query is progressively constructed. Detailed prompt templates are provided in Appendix LABEL:prompts.
@@ 4.1 Data Collection
To support queries that integrate both tabular and textual information, we augment the instance by leveraging the associated Wikipedia document. Specifically, certain entities within the query are replaced with reference-based formulations by the human annotators. For example, the query "Which driver " can be rephrased as "What is the nationality of the driver ". This entity substitution can either modify the subject of the question and its corresponding answer or alter the query phrasing while preserving the original answer. The annotation guidelines and annotator profiles are detailed in Appendix LABEL:instructions.
@@ 5.1.2 Implementation Details
In the text retrieval process, we employ the BGE-M3 series models (Chen et al., 2024a, b). During recall, we retain the top 30 candidates, from which the top 3 are subsequently selected via reranking. To manage large inputs, the text is chunked into segments of 1000 tokens, with a 200-token overlap between consecutive chunks. The iterative loop is bounded by a maximum of 5 iterations. For backbone LLMs, we utilize Claude-3.5-Sonnet as a representative closed-source LLM, while Deepseek-V3, Deepseek-R1 (Guo et al., 2025), and Qwen-2.5-72B (Yang et al., 2024) serve as the open-source counterparts. A consistent backend is maintained for all modules in the online iterative reasoning process. We use accuracy as the evaluation metric, assessed by Qwen-2.5-72B, which yields a binary score of 0 or 1. The prompt is provided in Appendix LABEL:prompts, and the results evaluated using the exact match metric are presented in Appendix LABEL:append_exp.
@@ 5.1.3 Baselines
We evaluate the performance of TableRAG by benchmarking it against three distinct baseline methodologies: (1) Direct answer generation with LLMs. (2) NaiveRAG, which processes tabular data as linearized Markdown formatted texts and subsequently applies a standard RAG pipeline. (3) ReAct (Yao et al., 2023), a prompt based paradigm to synergize reasoning and acting in LLMs with external knowledge sources. (4) TableGPT2 (Su et al., 2024) employs a Python-based execution module to generate code (e.g., Pandas) for answer derivation within a simulated environment. The detailed implementation of these baseline methods is provided in Appendix LABEL:baseline and additional baseline comparisons are presented in Appendix LABEL:append_exp.
@@ 5.2 Main Result
The main results across different LLMs backbones are presented in Table 1. Several key observations emerge: (1) The ReAct framework demonstrates advantages over naive RAG on multi-source data, but exhibits degraded performance on single-source data that requires tabular reasoning. This can be attributed to context insensitivity during multi-turn reasoning. Queries that could be resolved with single SQL execution are instead decomposed into multiple subqueries. This over-fragmentation may introduce errors or incomplete information during execution (especially when operations such as filtering, aggregation are involved), potentially leading to cascading failures. (2) TableGPT2 yields acceptable results solely on single-source queries, such as WikiTQ, underscoring its limited capacity for handling multi-source queries. This reflects a lack of generalizability in heterogeneous information environments. (3) TableRAG surpasses all baselines, achieving at least a 10% improvement over the strongest alternative. Notably, it performs robustly across both single-source and multi-source data. This performance gain is attributed to the incorporation of symbolic reasoning, which enables effective adaptation to heterogeneous documents. Moreover, the consistency in performance across different LLM backbones underscores TableRAG’s architectural generality and compatibility with a broad range of backbones.
| Method | WikiTQ | HeteQA | ||
| Total Latency | Avg. Latency / Step | Total Latency | Avg. Latency / Step | |
| ReAct | 13.50 | 5.92 | 24.57 | 7.70 |
| TableRAG | 17.70 | 7.93 | 45.65 | 10.79 |
@@ 6.1 Execution Dynamics
6.2 Latency
We assess the latency performance using Qwen-72B-Instruct as a consistent backbone model. As shown in Table 2, TableRAG incurs lower latency compared to ReAct, in both average time per step and total time efficiency. This is mainly due to: (1) TableRAG executes fewer iterations, and employs symbolic SQL execution, which is computationally efficient. (2) ReAct performs reasoning at every step, and consumes more steps before reaching a final decision, thereby incurring greater time overhead.
@@ 7 Analysis
We provide a comprehensive analysis of TableRAG in this section, with additional results presented in Appendix LABEL:analysis.
@@ Appendix A HeteQA
| sql |